Assume that (G,*) is a group, that g is an element of G and that t is the smallest positive integer such that g^t=e

prove that g^n=e if and only if t |n. use division algorithm and explain why r must = 0
1
Expert's answer
2020-10-22T17:55:08-0400

Let us prove the converse first.

If "t|n" . thus there exist "k\in \Bbb Z" such that "kt=n" .

Now,

"g^n=g^{kt}=(g^t)^k=e^k=e"

Thus,

"g^n=e"

Suppose, "t" not divides "n" , hence by division algorithm there are p,r such that


"n=pt+r,\hspace{1cm}0<r<t"

Moreover we have given

"g^n=e\\
\implies g^n=g^{pt+r}\\
\implies(g^t)^pg^r\\
\implies e^pg^r\\
\implies eg^r=g^r=e"

But we have given that "t" is given smallest , hence by definition "t|r\implies r\ge t" , hence we get a contradiction.

Therefore r must 0, hence

"n=pt\implies t|n"

We are done.